OpenStudy (anonymous):

Solve the system by substitution or elimination. 4x – 6y = –1 8x + 3y = 3

OpenStudy (anonymous):

Let's do this one together. Ready?

OpenStudy (anonymous):

Would you like to cancel out the x's first or the y's first?

OpenStudy (anonymous):

X'S FIRST

OpenStudy (anonymous):

Awesome. So we have 4x and 8x to work with. Your best bet is to change the 4x into -8x. What would we multiply 4x by to get -8x?

OpenStudy (anonymous):

Also, let me know if it makes sense why we're trying to get -8x. If you don't understand, please tell me and I'll explain.

OpenStudy (anonymous):

EXPLAIN MORE PLEASE KINDA CONFUSED

OpenStudy (anonymous):

Sure thing! So you see that our original equations were 4x – 6y = –1 8x + 3y = 3

OpenStudy (anonymous):

Now, if we were to add them together like they are right now, we'd have 12x -3y = 2 However, we want to multiply one of the equations by something that would give us 0x.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

So, if we keep our 8x + 3y = 3 just the way it is, we'd want equation 2 to say -8x somewhere. That way, you'd have 8x - 8x + (blah blah) = blah blah. The 8x and the -8x would kill the x's completely so that all we have to work with is y. So, what do we multiply the equation, 4x – 6y = –1 by to get -8x in the front?

OpenStudy (anonymous):

yes I understand that now, now what about the other part?

OpenStudy (anonymous):

Well, tell me what number we'll multiply by so that we can do the next step.

OpenStudy (anonymous):

would it be the 3?

OpenStudy (anonymous):

If you multiply by 3, you'd get this: 3(4x – 6y = –1) 8x + 3y = 3 Which would turn into 12x -18y = -3 8x + 3y = 3 And adding the x's would give you 20x, not 0x. Remember, you want to multiply the 4x by something to give you -8x, so that when you add the -8x to 8x, you get 0x. So what do we multiply 4x by to get 8-x?

OpenStudy (anonymous):

I meant "...4x by to get -8x?"

OpenStudy (anonymous):

3?